OpenStudy (anonymous):

You are paying $5 to get in the matinee plus $2 for each concession item you buy. You buy x items at the concessions. Write a mathematical expression for the amount of money you are spending at the movies

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

What are you being asked for? @reyna24

OpenStudy (anonymous):

does that help?

OpenStudy (skullpatrol):

@aysian97 that is the answer

OpenStudy (anonymous):

@reyna24, do you want me to explain it to you?

OpenStudy (anonymous):

thank you what about this one The perimeter of a rectangle is the sum of all of its sides. Write a mathematical expression for the perimeter of a rectangle with a length of 4 and a width of w. Simplify the expression if possible

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

sorry, @skullpatrol

OpenStudy (anonymous):

@reyna24, would you like me to show you how i got that?

OpenStudy (anonymous):

okay, so your trying to find the perimeter. In order to find the perimeter of a rectangle, you have to add up all the sides. Now you know that the length is 4, and if you have TWO side lengths then you have to multiply that by 2. so it would be 8. Now, you have the width. The question does not give you the width, so you just put a variable, "w". Sooo, now you have "w" as the width. You have two widths, so the expression would look like, "2w". Now, you put it all together. p=8+2w. You don't have the value of "w", or the value of "p". Therefore, you can't go on. Does that make sense?

OpenStudy (anonymous):

what about this one give an example of the commutative property of addition

OpenStudy (anonymous):

okay, can you tell me what the commutative property of addition is? In your own words...

OpenStudy (anonymous):

seee thing i dont no

OpenStudy (anonymous):

okay let me give you the definition...and then you try and make an equation out of it.

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

okay here's the definition: The Commutative Property of Addition states that changing the order of addends does not change the sum, i.e. if a and b are two real numbers, then a + b = b + a.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

your welcome :) glad to help